# Transfer learning on multi-dimensional data: a novel approach to neural network-based surrogate modeling

### v1

@@ Abstract
The development of efficient surrogates of partial differential equations (PDEs) is a critical step towards scalable modeling of complex, multiscale systems-of-systems. Convolutional neural networks (CNNs) have gained popularity as the basis for such surrogate models due to their success in capturing high-dimensional input-output mappings and the negligible cost of a forward pass. However, the high cost of generating training data – typically via classical numerical solvers – raises the question of whether these models are worth pursuing over more straightforward alternatives with well-established theoretical foundations, such as Monte Carlo methods. To reduce the cost of data generation, we propose training a CNN surrogate model on a mixture of numerical solutions to both the $d$-dimensional problem and its ($d-1$)-dimensional approximation, taking advantage of the efficiency savings guaranteed by the curse of dimensionality. We demonstrate our approach on a multiphase flow test problem, using transfer learning to train a dense fully-convolutional encoder-decoder CNN on the two classes of data. Numerical results from a sample uncertainty quantification task demonstrate that our surrogate model outperforms Monte Carlo with several times the data generation budget.
@@ 1 Introduction
A number of strategies for surrogate modeling have been proposed – for example, Kriging and Gaussian processes [2, 3, 4]; polynomial chaos expansion [5, 6]; support vector regression [7]; and, recently, deep neural networks [e.g. 8]. Surrogates based on neural networks (NNs), and convolutional neural networks (CNNs) in particular, are a natural choice for this task due to their ability to handle high-dimensional input-output mappings [9] and the negligible cost of a single forward pass. CNNs have proven particularly successful in image processing tasks [10, 11], a quality which we and others leverage by casting the surrogate modeling problem as an image-to-image regression problem.
@@
To see this, we will consider a typical setting for UQ. Suppose, as is common in many scientific applications, that we wish to analyze some output variable of a PDE on domain $\Omega$, whose true value is denoted by $Q:\Omega\rightarrow\mathbb{R}$, subject to a random input variable $\xi$ such that $Q=Q(\xi)$. [^2] This extends to sets of random variables as well as random functions on the domain. In particular, we aim to estimate $\mathbb{E}_{\xi}[Q]$. [^3] Our analysis extends readily to other moments, and even full PDFs and CDFs. Let $Q_{M}=Q_{M}(X_{M})\in\mathbb{R}$ denote the approximation of $Q$ from the numerical solution to the PDE [^4] For example, via finite difference, finite element, or finite volume methods on a discretized computational domain of $M$ nodes, where $X_{M}:=(\xi_{1},...,\xi_{M})$ is the vector of random inputs on these nodes. We assume that $\mathbb{E}[Q_{M}]\rightarrow\mathbb{E}[Q]$ as $M\rightarrow\infty$.
Given a fixed data generation budget (for example, sufficient to run our PDE solver $N$ times, on $N$ samples of $X_{M}$), two possible approaches to estimating $\mathbb{E}_{\xi}[Q]$ are to construct a typical Monte Carlo estimator $\hat{Q}_{MC}$ with the $N$ solutions, or to train a NN-based surrogate model on the $N$ solutions and use the surrogate to estimate $\hat{Q}_{NN}$. We’ll denote by $\hat{Q}_{M}$ our chosen estimator of $\mathbb{E}[Q_{M}]$, and quantify the accuracy of this estimator via its mean square error (MSE) (also called the average discrepancy):
@@
Therefore, for a suitably refined discretization, the average discrepany of an MC estimator for $\mathbb{E}[Q]$ decays inversely with the number of samples – a well-known result. In contrast, our NN-based surrogate model estimator $\hat{Q}_{NN}$ cannot be assumed to be unbiased. We therefore have:
@@
where $\epsilon^{2}_{*}$ cannot be further specified without assumptions on the bias of the NN estimator and on the correlation between the discretization error and estimator error.
@@ 2 Methods
A fundamental assumption underlying multifidelity learning is that the low-fidelity data, while less accurate than the high-fidelity data, still contains valuable information that can help the model learn the desired input-output mapping. A similar intuition underlies transfer learning, which operates on the premise that the learning performed for one modeling task is still valuable when applied to a new task. Alternative approaches to multifidelity learning, for example explicitly estimating the correlation between high- and low-fidelity data and embedding this into the model [21], have also proven succesful. In the present work, the natural alignment between multifidelity learning and transfer learning motivates our use of transfer learning to implement our multifidelity approach.
In the following sections, we briefly describe our model architecture and transfer learning strategy, as well as the computational test problem, data generation process, and computational resources we used to generate our results.
@@
We take a similar approach to [17], adapting the DenseED architecture proposed by [22] to accommodate a transfer learning strategy. DenseED is a dense fully convolutional encoder-decoder CNN which has been successfully used to develop surrogate models for UQ in groundwater [23] and multiphase flow [24] problems.
@@
Specifically, we first train a low-fidelity model using the $(d-1)$-dimensional data, then use the resulting model to intialize a high-fidelity model trained on the $d$-dimensional data. This provides a straightforward means of incorporating both sets of training data, such that the bulk of the training is performed on the more abundant low-fidelity data, while the high-fidelity data is used to refine the model predictions.
We implement this transfer learning strategy by training in three phases, following the work of [17]. In Phase 1, we train a low-fidelity CNN surrogate exclusively on $(d-1)$-dimensional data. In Phase 2, we freeze all CNN weights excluding those in the last convolutional layer, and retrain this last layer using the $d$-dimensional data. Finally, in Phase 3, we unfreeze all weights, and retrain the entire network with the $d$-dimensional data. Since the weights in both Phase 2 and Phase 3 are initialized with the weights from the previous phase, this transfer learning strategy ultimately yields a high-fidelity surrogate model with far fewer high-fidelity training samples than would otherwise be required.
@@
To demonstrate our proposed method, we consider a multiphase Darcy flow problem for two incompressible and immiscible fluids, denoted by $l\in\{1,2\}$. This problem is pervasive in science and engineering applications, and particularly relevant for applications such as carbon sequestration. The governing equations are the coupled PDEs relating saturation $S_{l}(\mathbf{x},t)$, velocity $\mathbf{v}_{l}(\mathbf{x},t)$, and pressure $P_{l}(\mathbf{x},t)$:
@@
We consider a 2D, square domain $\Omega$ of size $150\times 150$ with Dirichlet boundary conditions on the left and right boundaries, and no-flow conditions on the upper and lower boundaries:
@@
Since our model architecture is designed to map 2D inputs to 2D outputs, we scale up the low-fidelity data (both inputs and outputs) to match the dimensions of the high-fidelity data – in this case, $n\times n$. We test two alternative methods to do this, which we refer to as the “high-frequency” and “low-frequency” approaches. In the high-frequency approach, we concatenate $n$ separate 1D permeability fields to generate a single input image and $n$ corresponding 1D solutions to generate the corresponding output image. In the low-frequency approach, we replicate a single 1D permeability field $n$ times to generate an input image, and replicate the corresponding 1D solution $n$ times to generate the output image. An example of an input-output image pair for the low frequency approach is provided in Figure 1.
@@ 3 Results
We tested our CNN surrogate model in an uncertainty quantification task, comparing its performance to that of a standard Monte Carlo approach. The task at hand was to constrain the PDF of the breakthrough time, $T_{b}$, where
@@
corresponding to the first time at which the invading phase reaches a desired saturation threshold (in this case, 0.15) at a given point in the domain (here, $x_{1}=100$). Constraining the breakthrough time of an invading fluid is a common problem in hydrology and reservoir engineering, for example, to predict when pollutants may reach drinking water sources [30], or when injected fluid will reach production wells. We approximated the PDF of $T_{b}$ by evaluating 2000 forward passes of our multifidelty CNN surrogate. This process took less than 18 seconds to complete once the model was trained. We compared the output to the PDF obtained from Monte Carlo with as many forward passes of the original high-fidelity numerical solver as possible for a given data generation budget. We found that the CNN surrogate, trained on only 4 hours’ worth of multifidelity data, performed comparably to Monte Carlo on 24-48 hours’ worth of high-fidelity numerical solutions. This result was robust to multiple error metrics, as depicted in Figure 2.
Figure 2: UQ results: mean average error (MAE) and Kullback-Leibler (KL) divergence of the PDF of $T_{b}$. From left: 2000 forward passes of the CNN surrogate trained on 4 hours’ worth of multi-dimensional data (in purple); standard Monte Carlo with 6, 12, 24, and 48 hours’ worth of high fidelity runs (in progressively darker green). For both measures, the error of the CNN surrogate is nearly as low as or lower than that of Monte Carlo conducted on 24 hours’ worth of high fidelity runs (and with less variance). Error bars represent standard deviation over 50 trials.
@@
We found that balancing the training data generation budget between low- and high-fidelity data was optimal from the perspective of minimizing the root mean square error (RMSE). At the two extremes of: 1) nearly all high-fidelity data (and relatively few training samples); and 2) nearly all low-fidelity data (and relatively many training samples), the models performed worse than when the data generation budget was more evenly split. This result is illustrated in Figure 3, where the x-axis indicates an increasing number of 2D samples (bottom axis) and decreasing number of 1D samples (top axis) from left to right. This result was also consistent across both the high- and low-frequency approaches to generating low-fidelity data. These results are consistent with prior work [17], and suggest that balancing 1D and 2D data in the training set allows the model to benefit from both the accuracy of the high-fidelity 2D data and the larger sample size afforded by the low-fidelity 1D data.
@@
We found that the low-frequency approach to data generation significantly outperformed the high-frequency approach, despite containing less information and taking a fraction of the time to generate (just over 1 second versus 133 seconds). In Figure 3, the purple and blue curves show the high- and low-frequency results, respectively, each trained with the same number of high- and low-fidelity training samples. [^7] For the low-frequency results, this means that data-generation did not utilize the full allotted budget of 4 hours. This behavior is consistent with the well-known spectral bias of neural networks [31, 32]. We also found that increasing the number of low-frequency low-fidelity training samples to fully utilize the 4 hour data-generation budget further improved the performance of the surrogate. In fact, for several different splits between 1D and 2D data, our multifidelity model achieved lower error than even the CNN surrogate trained on 12 hours’ worth of 2D data. This finding is shown by the green curve in Figure 3. These results are encouraging as they indicate that supplementing a limited data generation budget with inexpensive ($d-1$)-dimensional data can lead to substantial improvements in model performance.
@@ 4 Discussion
We demonstrated that a CNN surrogate model for a 2D PDE can be trained on a mixture of numerical solutions to the 2D PDE and numerical solutions to its 1D approximation. This approach allows us to perform UQ tasks with lower error compared to Monte Carlo using several times the data generation budget. These results suggest that the curse of dimensionality may be tamed and even exploited in our favor; the ability of CNNs to learn useful features from a lower-dimensional approximation to the original problem allows us to generate much more training data than would otherwise be possible, given a fixed data generation budget and the typically high cost of running physics-based simulations.
@@
We note that the results described here were achieved with a relatively off-the-shelf method. We selected the current model architecture based on its success in generating a surrogate model for similar problems [17, 22, 23, 24], but acknowledge that several measures could be taken to improve its performance. For example, we did not conduct any hyper-parameter optimization, and rather adopted the parameters identified by [17]. Data augmentation (e.g. adding rotated or flipped versions of training images to the training dataset as if they were independent samples) has been shown to improve NN performance [33]. This technique could potentially improve the ability of the ($d-1$)-dimensional dataset to capture dynamics in the missing dimension, or augment a small $d$-dimensional dataset. In terms of architecture, other studies on multifidelity learning emphasize the benefits of computing and exploiting the correlation between high- and low-fidelity data [18, 19, 21]. Implementing memory-aware synapses, as in [34], could also ensure that the weights of the optimal layer are being retrained in each phase, so that useful weights are maintained and less useful weights are refined. Finally, we noted the occurrence of aliasing in the output images. While this is not a major issue for tasks like uncertainty quantification, techniques such as Fourier feature embedding [35] can help address this phenomenon, further improving performance and accuracy in capturing the sharp saturation front.
@@
Our results suggest that it is possible to tame the curse of dimensionality by training a CNN-based surrogate model on multi-dimensional data. Our approach harnesses the strength of CNNs in capturing spatial relationships across multiple dimensions, while overcoming the slow convergence of NN-based surrogates with respect to training set size – one of the fundamental challenges with NN-based prediction. While further study is needed to verify this approach on higher-dimensional problems and additional classes of PDEs, we expect that our results should readily extend to these regimes. The continued development of simple-yet-effective approaches such as the one proposed here could significantly broaden the applicability of CNN-based surrogate models, providing powerful tools for researchers dealing with high-dimensional UQ tasks across various scientific and engineering domains.

### v2

@@ Abstract
The development of efficient surrogates for partial differential equations (PDEs) is a critical step towards scalable modeling of complex, multiscale systems-of-systems. Convolutional neural networks (CNNs) have gained popularity as the basis for such surrogate models due to their success in capturing high-dimensional input-output mappings and the negligible cost of a forward pass. However, the high cost of generating training data — typically via classical numerical solvers — raises the question of whether these models are worth pursuing over more straightforward alternatives with well-established theoretical foundations, such as Monte Carlo methods. To reduce the cost of data generation, we propose training a CNN surrogate model on a mixture of numerical solutions to both the $d$-dimensional problem and its ($d-1$)-dimensional approximation, taking advantage of the efficiency savings guaranteed by the curse of dimensionality. We demonstrate our approach on a multiphase flow test problem, using transfer learning to train a dense fully-convolutional encoder-decoder CNN on the two classes of data. Numerical results from a sample uncertainty quantification task demonstrate that our surrogate model outperforms Monte Carlo with several times the data generation budget.
@@ 1 Introduction
A number of strategies for surrogate modeling have been proposed — for example, Kriging and Gaussian processes [2, 3, 4]; polynomial chaos expansion [5, 6]; support vector regression [7]; and, recently, deep neural networks [e.g., 8]. Surrogates based on neural networks (NNs), and convolutional neural networks (CNNs) in particular, are a natural choice for this task due to their ability to handle high-dimensional input-output mappings [9] and the negligible cost of a single forward pass. CNNs have proven particularly successful in image processing tasks [10, 11], a quality which we and others leverage by casting the surrogate modeling problem as an image-to-image regression problem.
@@
To illustrate this, we will consider a typical setting for UQ. Suppose, as is common in many scientific applications, that we wish to analyze some output variable of a PDE on domain $\Omega$, whose true value is denoted by $Q:\Omega\rightarrow\mathbb{R}$, subject to a random input variable $\xi$ such that $Q=Q(\xi)$. [^2] This extends to sets of random variables as well as random functions on the domain. In particular, we aim to estimate $\mathbb{E}_{\xi}[Q]$. [^3] Our analysis extends readily to other moments, and even full PDFs and CDFs. Let $Q_{M}=Q_{M}(X_{M})\in\mathbb{R}$ denote the approximation of $Q$ from the numerical solution to the PDE [^4] For example, via finite difference, finite element, or finite volume methods on a discretized computational domain of $M$ nodes, where $X_{M}:=(\xi_{1},...,\xi_{M})$ is the vector of random inputs on these nodes. We assume that $\mathbb{E}[Q_{M}]\rightarrow\mathbb{E}[Q]$ as $M\rightarrow\infty$.
Given a fixed data generation budget (for example, sufficient to run our PDE solver $N$ times, on $N$ samples of $X_{M}$), two possible approaches to estimating $\mathbb{E}_{\xi}[Q]$ are to construct a typical Monte Carlo estimator $\hat{Q}_{MC}$ with the $N$ solutions, or to train a NN-based surrogate model on the $N$ solutions and use the surrogate to estimate $\hat{Q}_{NN}$. We denote our chosen estimator of $\mathbb{E}[Q_{M}]$ by $\hat{Q}_{M}$ and quantify the accuracy of this estimator via its mean square error (MSE) (also called the average discrepancy):
@@
Therefore, for a suitably refined discretization, the average discrepany of an MC estimator for $\mathbb{E}[Q]$ decays inversely with the number of samples — a well-known result. In contrast, our NN-based surrogate model estimator $\hat{Q}_{NN}$ cannot be assumed to be unbiased. We therefore have:
@@
where $\epsilon^{2}_{*}$ cannot be further specified without assumptions on the bias of the NN estimator and the correlation between the discretization error and estimator error.
@@ 2 Methods
A fundamental assumption underlying multifidelity learning is that the low-fidelity data, while less accurate than the high-fidelity data, still contains valuable information that can help the model learn the desired input-output mapping. A similar intuition underlies transfer learning, which operates on the premise that the learning performed for one modeling task is still valuable when applied to a new task. Alternative approaches to multifidelity learning, for example explicitly estimating the correlation between high- and low-fidelity data and embedding this into the model [21], have also proven successful. In the present work, the natural alignment between multifidelity learning and transfer learning motivates our use of transfer learning to implement our multifidelity approach.
In the remainder of this section, we briefly describe our model architecture and transfer learning strategy, as well as the computational test problem, data generation process, and computational resources we used to generate our results.
@@
We take a similar approach to [17], adapting the DenseED architecture proposed by [22] to accommodate a transfer learning strategy. DenseED is a dense fully convolutional encoder-decoder CNN that has been successfully used to develop surrogate models for UQ in groundwater [23] and multiphase flow [24] problems.
@@
Specifically, we first train a low-fidelity model using the $(d-1)$-dimensional data, then use the resulting model to initialize a high-fidelity model trained on the $d$-dimensional data. This provides a straightforward means of incorporating both sets of training data, such that the bulk of the training is performed on the more abundant low-fidelity data, while the high-fidelity data is used to refine the model predictions.
We implement this transfer learning strategy by training in three phases, building on the work of [17]. In Phase 1, we train a low-fidelity CNN surrogate exclusively on $(d-1)$-dimensional data. In Phase 2, we freeze all CNN weights excluding those in the final convolutional layer, which we retrain using the $d$-dimensional data. Finally, in Phase 3, we unfreeze all weights and retrain the entire network with the $d$-dimensional data. By initializing the weights in Phases 2 and 3 with those from the previous phase, our training strategy enables the model to leverage earlier training and reduces the number of high-fidelity $d$-dimensional training samples ultimately required. . Figure 1 illustrates the three training phases, highlighting the respective inputs and outputs for the model at each stage.
Figure 1: We train our CNN surrogate in 3 phases. Phase 1 produces a low-fidelity surrogate trained exclusively on the 1D data, scaled up to 2D for compatibility. Phase 2 produces a high-fidelity surrogate in which all weights excluding those in the last layer are identical to those of the low-fidelity model, but the weights in the last layer are retrained with the 2D data. Phase 3 is initialized with the model weights from Phase 2, then all weights are unfrozen and retrained again with the 2D data, producing a refined high-fidelity surrogate model.
@@
To demonstrate our proposed method, we consider a multiphase Darcy flow problem for two incompressible and immiscible fluids, denoted by $l\in\{1,2\}$. This problem is pervasive in science and engineering applications and is particularly relevant for applications such as carbon sequestration. The governing equations are the coupled PDEs relating saturation $S_{l}(\mathbf{x},t)$, velocity $\mathbf{v}_{l}(\mathbf{x},t)$, and pressure $P_{l}(\mathbf{x},t)$:
@@
We consider a 2D, square domain $\Omega$ of size $150\times 150$ with Dirichlet boundary conditions on the left and right boundaries and no-flow conditions on the upper and lower boundaries:
@@
Since our model architecture is designed to map 2D inputs to 2D outputs, we scale up the low-fidelity data (both inputs and outputs) to match the dimensions of the high-fidelity data — in this case, $n\times n$. We test two alternative methods to do this, which we refer to as the “high-frequency” and “low-frequency” approaches. In the high-frequency approach, we concatenate $n$ separate 1D permeability fields to generate a single input image and $n$ corresponding 1D solutions to generate the corresponding output image. In the low-frequency approach, we replicate a single 1D permeability field $n$ times to generate an input image, and replicate the corresponding 1D solution $n$ times to generate the output image. An example of an input-output image pair for the low frequency approach is provided in Figure 2. Note that this design choice affects the training dataset for Phase 1 but not Phase 2 or Phase 3.
@@ 3 Results
In this section we present three major insights gleaned from our numerical experiments. The first pertains to the viability of the method: we find that our surrogate model outperforms standard Monte Carlo in uncertainty quantification tasks, even with less data. We also present two findings relevant to the generation and setup of the training dataset.
@@
To evaluate the effectiveness of our model, we applied it to a sample uncertainty quantification task. A common challenge in hydrology and reservoir engineering is predicting the breakthrough time of an invading fluid, which is critical for estimating when pollutants might reach drinking water sources or when injected fluids will arrive at production wells [30].
We let $T_{b}$ denote the breakthrough time:
@@
defined as the first time at which the invading phase reaches a desired saturation threshold (in this case, 0.15) at a given point in the domain (here, $x_{1}=100$). We approximated the PDF of $T_{b}$ by evaluating 2000 forward passes of our multifidelty CNN surrogate. This process took less than 18 seconds to complete once the model was trained. We compared the output to the PDF obtained from standard Monte Carlo with as many forward passes of the original high-fidelity numerical solver as possible for a given data generation budget. We found that the CNN surrogate, trained on only 4 hours’ worth of multifidelity data, performed comparably to Monte Carlo on 24-48 hours’ worth of high-fidelity numerical solutions. This result was robust to multiple error metrics, as depicted in Figure 3.
Figure 3: UQ results: mean average error (MAE) and Kullback-Leibler (KL) divergence of the PDF of $T_{b}$. From left: 2000 forward passes of the CNN surrogate trained on 4 hours of multi-dimensional data (in purple); standard Monte Carlo with 6, 12, 24, and 48 hours of high fidelity runs (in progressively darker green). For both measures, the CNN surrogate achieves similar or better results than the Monte Carlo method on 24 hours of high fidelity runs, with lower variance. Error bars represent the standard deviation over 50 trials.
@@
We found that balancing the training data generation budget between low- and high-fidelity data was optimal from the perspective of minimizing the root mean square error (RMSE). At the two extremes of: 1) nearly all high-fidelity data (and relatively few training samples); and 2) nearly all low-fidelity data (and relatively many training samples), the models performed worse than when the data generation budget was more evenly split. This result is illustrated in Figure 4, where the x-axis indicates an increasing number of 2D samples (bottom axis) and decreasing number of 1D samples (top axis) from left to right. This result was also consistent across both the high- and low-frequency approaches to generating low-fidelity data. This finding is consistent with prior work [17], and suggests that balancing 1D and 2D data in the training set allows the model to benefit from both the accuracy of the high-fidelity 2D data and the larger sample size afforded by the low-fidelity 1D data.
@@
We found that the low-frequency approach to data generation significantly outperformed the high-frequency approach, despite containing less information and taking a fraction of the time to generate (just over 1 second versus 133 seconds). This can be seen in Figure 4, where the purple and blue curves show the high- and low-frequency results, respectively, each trained with the same number of high- and low-fidelity training samples. [^7] For the low-frequency results, this means that data-generation did not utilize the full allotted budget of 4 hours. This behavior is consistent with the well-known spectral bias of neural networks [31, 32]. These results are also consistent with our expectations — given that our ultimate goal is to model a single flow scenario, the weights learned from modeling multiple flow scenarios may not be as relevant as those learned from modeling a single flow scenario.
We also found that increasing the number of low-frequency low-fidelity training samples to fully utilize the 4 hour data-generation budget further improved the performance of the surrogate. In fact, for several different splits between 1D and 2D data, our multifidelity model achieved lower error than even the CNN surrogate trained on 12 hours’ worth of 2D data. These findings are shown by the green curve in Figure 4. These results are encouraging as they indicate that supplementing a limited data generation budget with inexpensive ($d-1$)-dimensional data can lead to substantial improvements in model performance.
@@ 4 Discussion
In this work, we demonstrate a novel approach to training CNN-based surrogate models that addresses a fundamental but often-ignored challenge with such methods: slow convergence with respect to training dataset size. Specifically, we demonstrate that a CNN surrogate model for a 2D PDE can be trained on a mixture of numerical solutions to the 2D PDE and numerical solutions to its 1D approximation. With this approach, we achieve lower error in UQ tasks compared to Monte Carlo using several times the data generation budget. These results suggest that the curse of dimensionality may be tamed and even exploited in our favor; the ability of CNNs to learn useful features from a lower-dimensional approximation to the original problem allows us to generate much more training data than would otherwise be possible, given a fixed data generation budget and the typically high cost of running physics-based simulations.
@@
We note that the results described here are achieved with a relatively off-the-shelf method. We selected the current model architecture based on its success in generating a surrogate model for similar problems [17, 22, 23, 24], but acknowledge that several measures could be taken to improve its performance. For example, we did not conduct any hyper-parameter optimization, and rather adopted the parameters identified by [17]. Data augmentation (e.g., adding rotated or flipped versions of training images to the training dataset as if they were independent samples) has been shown to improve NN performance [33]. This technique could potentially improve the ability of the ($d-1$)-dimensional dataset to capture dynamics in the missing dimension, or augment a small $d$-dimensional dataset. In terms of architecture, other studies on multifidelity learning emphasize the benefits of computing and exploiting the correlation between high- and low-fidelity data [18, 19, 21]. Implementing memory-aware synapses, as in [34], could also ensure that the weights of the optimal layer are being retrained in each phase, so that useful weights are maintained and less useful weights are refined. Finally, we noted the occurrence of aliasing in the output images. While this is not a major issue for tasks like uncertainty quantification, techniques such as Fourier feature embedding [35] can help address this phenomenon, further improving performance and accuracy in capturing the sharp saturation front.
@@
Our results suggest that it may be possible to tame the curse of dimensionality by training a CNN-based surrogate model on multi-dimensional data. Our approach harnesses the strength of CNNs in capturing spatial relationships across multiple dimensions, while overcoming the slow convergence of NN-based surrogates with respect to training set size — one of the fundamental challenges with NN-based prediction. While further study is needed to verify this approach on higher-dimensional problems and additional classes of PDEs, we expect that our results should readily extend to these regimes. The continued development of simple-yet-effective approaches such as the one proposed here could significantly broaden the applicability of CNN-based surrogate models, providing powerful tools for researchers dealing with high-dimensional UQ tasks across various scientific and engineering domains.